# Leveraging Confident Image Regions for Source-Free Domain-Adaptive Object Detection

### v1

@@ Abstract
Source-free domain-adaptive object detection is an interesting but scarcely addressed topic. It aims at adapting a source-pretrained detector to a distinct target domain without resorting to source data during adaptation. So far, there is no data augmentation scheme tailored to source-free domain-adaptive object detection. To this end, this paper presents a novel data augmentation approach that cuts out target image regions where the detector is confident, augments them along with their respective pseudo-labels, and joins them into a challenging target image to adapt the detector. As the source data is out of reach during adaptation, we implement our approach within a teacher-student learning paradigm to ensure that the model does not collapse during the adaptation procedure. We evaluated our approach on three adaptation benchmarks of traffic scenes, scoring new state-of-the-art on two of them. The code will be released upon acceptance.
Index Terms:
Object detection, domain adaptation, cross-domain shift, data augmentation.I Introduction
@@ II-C Source-free unsupervised domain adaptation for object detection
Method | Setting | In a nutshell | Base detector | Source Target | Code |
ConfMix [30] | UDA | Self-training Cross-domain image mixing | YOLOv5 | C F, S C K C | link |
DACA [31] | UDA | Self-training Target-only region mixing Data augmentation | YOLOv5 | C F S C K C | link |
DA-Detect [24] | UDA | Image-level and object-level adaptation Adversarial mining for hard examples Auxiliary domain generation | Faster R-CNN | C F C K | link |
SC-UDA [52] | UDA | Intermediate domain generation with optimized CycleGAN Iterative self-training Uncertainty-based pseudo-label optimization | Faster R-CNN | C F S C K C | ✗ |
ViSGA [37] | UDA | Instance-level feature aggregation Similarity-based RPN proposal clustering Alignment via adversarial training | Faster R-CNN | C F S C K C | ✗ |
SOAP [50] | SF-UDA | Mean teacher Perturbed target images | Faster R-CNN | C F, S C K C, P W | link |
SFOD-Mosaic [26] | SF-UDA | Mining pseudo-label threshold via self-entropy descent | Faster R-CNN | C F, S C K C, C B | ✗ |
LODS [25] | SF-UDA | Mean teacher Non-linear target style enhancement | Faster R-CNN | C F, K C P Cl, P W | link |
ASFOD [5] | SF-UDA | Student-teacher Adversarial learning Source-similarity-based target division | Faster R-CNN | C F, S C K C, C B | link |
LPU [4] | SF-UDA | Mean teacher Confident pseudo-labels Proposal soft training on low-confident pseudo-detections | Faster R-CNN | C F, S C K C, C B | ✗ |
SF-UT [14] | SF-UDA | Mean teacher Adaptive Batch Normalization | Faster R-CNN | C F, S C K C | link |
SF-DACA (ours) | SF-UDA | Mean teacher Target-only region mixing Data augmentation | Faster R-CNN | C F S C K C | Soon |
@@ IV-C Results and discussion
1 | Source-only | 35.20 | 51.30 | 1.70 | 41.90 | 12.90 | 18.70 | 29.40 | 32.20 | 27.90 |
2 | Oracle | 48.60 | 67.10 | 25.80 | 52.70 | 35.10 | 38.80 | 44.90 | 53.70 | 45.90 |
3 | SOAP [50] | 35.90 | 48.40 | 24.30 | 45.00 | 23.90 | 31.80 | 37.90 | 37.20 | 35.50 |
4 | SFOD [26] | 33.20 | 44.50 | 22.20 | 40.70 | 25.50 | 28.40 | 34.10 | 39.00 | 33.50 |
5 | LODS [25] | 34.00 | 48.80 | 19.60 | 45.70 | 27.30 | 33.20 | 37.80 | 39.70 | 35.80 |
6 | ASFOD [5] | 32.30 | 44.60 | 29.00 | 44.10 | 28.10 | 31.80 | 38.90 | 34.30 | 35.40 |
7 | LPU [4] | 39.00 | 55.40 | 21.20 | 50.30 | 24.00 | 30.30 | 44.20 | 46.00 | 38.80 |
8 | SF-UT [14] | 40.90 | 58.90 | 50.20 | 48.00 | 29.60 | 36.20 | 44.10 | 51.90 | 45.00 |
9 | SF-DACA (ours) | 41.90 | 61.90 | 11.90 | 51.20 | 24.80 | 29.00 | 43.60 | 34.70 | 37.40 |
10 | Gain w.r.t. row 1 | +6.70 | +10.60 | +10.20 | +9.30 | +11.90 | +10.30 | +14.20 | +2.50 | +9.50 |
@@ IV-D Ablation study
1 | Without DACA | 34.20 | 55.30 | 58.30 |
2 | With DACA | 37.40 | 56.80 | 59.60 |
3 | Improvement | +3.20 | +1.50 | +1.30 |

### v2

@@ Abstract
Source-free domain-adaptive object detection is an interesting but scarcely addressed topic. It aims at adapting a source-pretrained detector to a distinct target domain without resorting to source data during adaptation. So far, there is no data augmentation scheme tailored to source-free domain-adaptive object detection. To this end, this paper presents a novel data augmentation approach that cuts out target image regions where the detector is confident, augments them along with their respective pseudo-labels, and joins them into a challenging target image to adapt the detector. As the source data is out of reach during adaptation, we implement our approach within a teacher-student learning paradigm to ensure that the model does not collapse during the adaptation procedure. We evaluated our approach on three adaptation benchmarks of traffic scenes, scoring new state-of-the-art on two of them.
I Introduction
@@ II-C Source-free unsupervised domain adaptation for object detection
| Method | Setting | In a nutshell | Base detector | Source Target | Code | ||||||
| ConfMix [30] | UDA |
YOLOv5 |
link | ||||||
| DACA [31] | UDA |
YOLOv5 |
link | ||||||
| DA-Detect [24] | UDA |
Faster R-CNN |
link | ||||||
| SC-UDA [52] | UDA |
Faster R-CNN |
| ViSGA [37] | UDA |
Faster R-CNN |
| SOAP [50] | SF-UDA |
Faster R-CNN |
link | ||||||
| SFOD-Mosaic [26] | SF-UDA |
Faster R-CNN |
| LODS [25] | SF-UDA |
Faster R-CNN |
link | ||||||
| ASFOD [5] | SF-UDA |
Faster R-CNN |
link | ||||||
| LPU [4] | SF-UDA |
Faster R-CNN |
| SF-UT [14] | SF-UDA |
Faster R-CNN |
link | ||||||
| SF-DACA (ours) | SF-UDA |
Faster R-CNN |
Soon |
@@ IV-C Results and discussion
| 1 | Source-only | 35.20 | 51.30 | 1.70 | 41.90 | 12.90 | 18.70 | 29.40 | 32.20 | 27.90 |
| 2 | Oracle | 48.60 | 67.10 | 25.80 | 52.70 | 35.10 | 38.80 | 44.90 | 53.70 | 45.90 |
| 3 | SOAP [50] | 35.90 | 48.40 | 24.30 | 45.00 | 23.90 | 31.80 | 37.90 | 37.20 | 35.50 |
| 4 | SFOD [26] | 33.20 | 44.50 | 22.20 | 40.70 | 25.50 | 28.40 | 34.10 | 39.00 | 33.50 |
| 5 | LODS [25] | 34.00 | 48.80 | 19.60 | 45.70 | 27.30 | 33.20 | 37.80 | 39.70 | 35.80 |
| 6 | ASFOD [5] | 32.30 | 44.60 | 29.00 | 44.10 | 28.10 | 31.80 | 38.90 | 34.30 | 35.40 |
| 7 | LPU [4] | 39.00 | 55.40 | 21.20 | 50.30 | 24.00 | 30.30 | 44.20 | 46.00 | 38.80 |
| 8 | SF-UT [14] | 40.90 | 58.90 | 50.20 | 48.00 | 29.60 | 36.20 | 44.10 | 51.90 | 45.00 |
| 9 | SF-DACA (ours) | 41.90 | 61.90 | 11.90 | 51.20 | 24.80 | 29.00 | 43.60 | 34.70 | 37.40 |
| 10 | Gain w.r.t. row 1 | +6.70 | +10.60 | +10.20 | +9.30 | +11.90 | +10.30 | +14.20 | +2.50 | +9.50 |
@@ IV-D Ablation study
| 1 | Without DACA | 34.20 | 55.30 | 58.30 |
| 2 | With DACA | 37.40 | 56.80 | 59.60 |
| 3 | Improvement | +3.20 | +1.50 | +1.30 |